COMPLEX DYNAMICS IN SEVERAL VARIABLES

JOHN SMILLIE GREGERY T. BUZZARD

Abstract. These notes are the outgrowth of a series of lectures given at MSRI in January 1995 at the beginning of the special semester in complex dynamics and hyperbolic geometry. In these notes, the primary aim is to motivate the study of complex dynamics in two variables and to introduce the major ideas in the field. Hence the treatment of the subject is mostly expository.